OpenStudy (anonymous):

more integration: (hopefully the last one)- (x^3)/(1+x^4)

OpenStudy (anonymous):

(1+x^4) = u du/4=x^3dx

OpenStudy (anonymous):

Take u=1+x^4, du=4(x^3)dx I think we can easily solve, yes?

OpenStudy (anonymous):

grrr I can never tell which method to use...I have all these tools in my tool box but can never figure out with one to use. thanks